# Good things come in small packages: Should we adopt Lite-GPUs in AI infrastructure?

### v1

@@ 5. Related Work
There is a plethora of work that propose systems solutions for improving performance (Agrawal et al., 2023; Dao et al., 2022), energy efficiency (Qiu et al., 2024; Stojkovic et al., 2024a), parallelism (Shoeybi et al., 2020; Liu et al., 2023b), and scheduling (Fu et al., 2024; Sun et al., 2024) of AI workloads in the data-center. These are complementary to the hardware and systems efforts on delivering cost-effective scaling of AI workloads using Lite-GPUs.
@@ 6. Conclusion
We are already facing uncertainty on the amount of compute and memory that can fit into a single GPU package, as cutting-edge GPUs already display the packaging, cooling, and cost-related challenges due to their complex designs. In this paper, we propose an alternative way of scaling AI infrastructure: by using Lite-GPUs instead of complex, expensive, and power-hungry large GPUs. Motivated by the yield, power, and operational benefits of smaller GPU packages, we look at AI infrastructure within the context of Lite-GPUs. We provide an overview of key research questions around workload, memory, and network management. We also present how Lite-GPUs can improve energy management, performance efficiency, and fault-tolerance. With this paper, we aim to start a discussion around Lite-GPUs and their potential to turn the tide on the many issues we face while building and operating GPU clusters in the era of generative AI.

### v2

@@ 5. Related Work
Alternatively to the scale-out approach of Lite-GPUs, wafer-scale computing systems aim to pack massive amounts of compute and communication bandwidth onto single, large integrated chips (Cerebras, [n. d.]; Hu et al., 2024). Though these systems benefit from much increased bandwidth and integration density, they require complex and advanced packaging techniques, which can lead to challenges around yield, cost, and power consumption (Hu et al., 2024).
There is a plethora of work that propose systems solutions for improving the performance (Agrawal et al., 2023; Dao et al., 2022), energy efficiency (Qiu et al., 2024; Stojkovic et al., 2024a), parallelism (Shoeybi et al., 2020; Liu et al., 2023b), and scheduling (Fu et al., 2024; Sun et al., 2024) of AI workloads in the data-center. Recently, DeepSeek demonstrated a variety of optimizations that enable efficient training and serving of a strong LLM on hardware that are relatively weaker than cutting-edge GPUs (DeepSeek-AI et al., 2025). These works are complementary to the hardware and systems efforts on delivering cost-effective scaling of AI workloads using Lite-GPUs.
@@ 6. Conclusion
We are already facing uncertainty on the amount of compute and memory that can fit into a single GPU package, as cutting-edge GPUs already display the packaging, cooling, power- and cost-related challenges of their complex designs. In this paper, we propose an alternative way of scaling AI infrastructure: by using Lite-GPUs instead of complex and expensive large GPUs. Motivated by the yield, power, and operational benefits of smaller GPU packages, we look at AI infrastructure within the context of Lite-GPUs. We provide an overview of key research questions around workload, memory, and network management. We also present how Lite-GPUs can improve energy management, performance efficiency, and fault-tolerance. With this paper, we aim to start a discussion around Lite-GPUs and their potential to turn the tide on the many issues we face while building and operating GPU clusters in the era of generative AI.